# Online Navigation Planning for Long-term Autonomous Operation of Underwater Gliders

### v1

@@ IV-D Simulator Parameter Tuning
To enable tuning $\Theta$, we assume availability of a dataset $\mathcal{D}=\{\mathbf{x}^{(i)}\}_{i=1}^{N}$ of historical dives with entries $\mathbf{x}^{(i)}$ of the form $((\mathbf{s}^{(i)},\mathbf{a}^{(i)},\beta^{(i)},\hat{\mathbf{q}}^{(i)},\mathbf{b}^{(i)}),\mathbf{s}_{\text{true}}^{(i)})$, comprising the conditions prior to starting a dive and the true post-dive state after the glider surfacing. For simplicity, we also assume that all parameters $\mathbf{c}$ except $n_{\text{yos}}$ and $z_{\text{bottom}}$ are held constant within a dataset used to fit the simulator parameters.
@@
We focus on optimising the generated simulator positions relative to the ground truth $\mathbf{p}^{(i)}_{\text{true}}$, as durations $\tau^{\prime}-\tau$ can realistically be treated as approximately constant given our assumption of fixed control parameters. For each $\mathbf{x}^{(i)}$, we draw $M$ samples of future states $\mathbf{s}^{(i,1)},\mathbf{s}^{(i,2)},\dots,\mathbf{s}^{(i,M)}\sim\mathcal{T}_{\Theta}(\mathbf{s},\mathbf{a},\beta,\hat{\mathbf{q}},\mathbf{b})$, and record the positions $\mathcal{P}^{(i)}=\{\mathbf{p}^{(i,1)},\mathbf{p}^{(i,2)},\dots,\mathbf{p}^{(i,M)}\}$. We subsequently fit a per-dive KDE with a Gaussian kernel $\hat{p}^{(i)}_{\Theta}=KDE(\mathcal{P}^{(i)};h)$ with bandwidth parameter $h$. Evaluating the probability $\hat{p}_{\Theta}^{(i)}(\mathbf{p}^{(i)}_{\text{true}})$ of the true surfacing location under the KDE gives us a score quantifying how well the simulator parameters $\Theta$ fit a particular dive. Furthermore, we also penalise the spread $\sum_{j=1}^{M}{(\mathbf{p}^{(i,j)}-\overline{\mathbf{p}^{(i,j)}})}^{2}$ of the positions relative to their mean. The scoring function $J(\Theta)$ is therefore defined as:
@@
$J(\Theta)=\sum_{i=1}^{N}{\log\hat{p}^{(i)}_{\Theta}(\mathbf{p}^{(i)}_{\text{true}})}-\lambda_{\text{reg}}\sum_{i=1}^{N}{\sum_{j=1}^{M}{(\mathbf{p}^{(i,j)}-\overline{\mathbf{p}^{(i,j)}})^{2}}},$
@@
where the first term corresponds to the log-likelihood $\ell(\Theta;\mathcal{D})$ and the second term penalises the spread of samples generated for each $\mathbf{x}^{(i)}$ with a configurable weight $\lambda_{\text{reg}}$.
@@
We seek to find the optimal parameters $\Theta^{*}=\operatorname*{\arg\!\max}_{\Theta}(J(\Theta))$. This is a high-dimensional optimisation problem with a scoring function that is expensive to evaluate and cannot be computed in closed form. As such, Bayesian optimization is a suitable framework. Concretely, we opt for BoTorch [32, 33], whose acquisition function navigates the exploration-exploitation tradeoff to generate promising candidate parameter sets $\Theta_{k}$ at each iteration $k$ of a total $n_{\text{SIM-BO}}$ iterations.
IV-E Sample-Based Online Planner
Gliders surface every few hours and await instructions for the next dive while uploading scientific data. Thus, an online planning approach is natural for solving this MDP, leveraging this time to derive a navigation plan. At each surfacing, the problem instance $I=(\mathbf{p}_{0},\tau_{0},\mathbf{p}_{\text{goal}},\rho,\mathbf{q}(\cdot),\mathbf{b}(\cdot))$ is assembled using the position $\mathbf{p}_{0}$ from the glider’s GPS fix, time $\tau_{0}$, goal specification, current forecasts, and bathymetry.
The online setup coupled with the sample-based nature of the simulator $\mathcal{T}_{\Theta}$ means UCT with double progressive widening is an appropriate choice. We note that, despite the discretisation of the actions, we opt for progressive widening in the action space also. This is done to maintain the branching factor low and increase the tree depth, enabling us to plan over a longer horizon.
Planning is carried out while the glider is at the surface using $n_{\text{trials}}$ samples, a parameter that is set on a per-deployment basis to account for the expected time duration of the glider at the surface and the computational resources available. As UCT is sequential, we opt for the root-parallel variant to aggregate results from $n_{\text{threads}}$ runs, improving the robustness of the navigation plan to noise in the current forecasts and action choices early in the tree search. Pseudocode for the algorithm is given in Algorithm 2. We highlight the following differences from a standard implementation:
  • Use of heuristic. We do not perform rollouts to estimate the value of leaf nodes as transitions require running a physics-based simulation. Instead, we use a heuristic (line 19) that estimates the time to reach the goal based on the current distance and constant speed. The error term $\epsilon_{\text{heur}}$ of the heuristic was calibrated based on preliminary deployment data and was set to approximately $1.77$. Equivalently, the typical trajectory length was $1.77\times$ that of the straight line.
@@ IV-E Sample-Based Online Planner
7:  end for
8:  $\mathbf{a}\leftarrow$ MajorityVote(allActions)
9:  return $\mathbf{a}$
10:Function RunUCT(II, randomState):
11:  $(\mathbf{p}_{0},\tau_{0},\mathbf{p}_{\text{goal}},\rho,\mathbf{q}(\cdot),\mathbf{b}(\cdot))\leftarrow\ I$
12:  create root node $v_{\text{root}}$ from $s_{0}=(\mathbf{p}_{0},\tau_{0})$
13:  for $i$ in $1,n_{\text{trials}}]$ do
14:   $v_{\text{leaf}}\leftarrow$ TreePolicy($v_{\text{root}}$)
15:   $\mathbf{s}\leftarrow$ GetState($v_{\text{leaf}}$)
16:   if $\mathbf{s}\in\mathcal{G}$ then
17:    $r\leftarrow$ SumCosts($\mathbf{s}_{0},\mathbf{s}$)
18:   else
19:    $r\leftarrow\epsilon_{\text{heur}}*(d(\mathbf{p},\mathbf{p}_{\text{goal}})/s)$
20:   end if
21:   Backup($v_{\text{leaf}},r$)
22:  end for
23:  action $\leftarrow$ RobustChild($v_{\text{root}}$)
24:  return action
V Glider Long-term Autonomy Engine
@@
In this section, we describe the components required for translating our methodological and algorithmic contributions discussed in the previous section to a field-validated system that performs long-term autonomous navigation planning for underwater gliders. The Glider Long-term Autonomy Engine system was designed and implemented in collaboration with glider pilots and marine robotics stakeholders in order to promote operational practicality and safety. The system as described herein is currently compatible with Slocum gliders [34, 35] and is in principle extensible to other designs such as Seaglider [36].
V-A Automated Goal Management
@@
The previous section specified problem instances $I$ with a single goal $\mathbf{p}_{\text{goal}}$. However, glider missions typically consist of many such goals. A common use case, which we target, is to have the glider sample data between a set of predefined points, i.e., a transect in the case of two points and the corners a polygon more generally. We therefore assume we are given an ordered list of points $\{p^{(i)}_{\text{goal}}\}_{i=1}^{n_{\text{goals}}}$. The system advances to the next goal in the list upon achieving the present one, and cycles to the beginning of the list once the end is reached.
V-B Waypoint List Generation
A waypoint list specifies an ordered sequence of waypoints that the glider should achieve. Waypoint lists are the first file type required for translating an action $\mathbf{a}=(\alpha,\mathbf{c})$ output by the planner to files that can be understood by a Slocum glider. While the action space of the planner is formulated in terms of relative bearings (Section IV-B), the Slocum glider expects a list of waypoints. A suitable translation approach is therefore needed.
Algorithm 3 Converting the output of the online planning algorithm to a list of waypoints as required by Slocum gliders.
1:Function ToWptList($\alpha,\mathbf{p}_{0},\mathbf{p}_{\text{goal}}^{\text{curr}},\mathbf{p}_{\text{goal}}^{\text{next}},\rho,\rho_{\text{wpt}},n_{\text{bck}}$):
2:  $\texttt{wpts}\leftarrow[\ ]$
3:  $\texttt{goal\_d}\leftarrow d(\mathbf{p}_{0},\mathbf{p}_{\text{goal}}^{\text{curr}})$
4:  if $\texttt{goal\_d}<\rho_{\text{wpt}}/2$ then
5:   wpts.append($\mathbf{p}_{\text{goal}}^{\text{curr}}$)
6:  else
7:   wpt_d $\leftarrow$ min(goal_d, $\rho_{\text{wpt}}$)
8:   $\beta$ = getBearing($\mathbf{p}_{0},\mathbf{p}_{\text{goal}}^{\text{curr}}$)
9:   $\psi\leftarrow\beta-\alpha$
10:   wpt $\leftarrow$ pointOnHeading($\mathbf{p}_{0},\psi$, wpt_d)
11:   wpts.append(wpt)
12:  end if
13:  addBackupWpts($\texttt{wpts},\mathbf{p}_{\text{goal}}^{\text{curr}},\rho,\rho_{\text{wpt}},n_{\text{bck}}$)
14:  addBackupWpts($\texttt{wpts},\mathbf{p}_{\text{goal}}^{\text{next}},\rho,\rho_{\text{wpt}},n_{\text{bck}}$)
15:Function addBackupWpts($\texttt{wpts},\mathbf{p}_{\text{goal}}^{\text{bck}},\rho,\rho_{\text{wpt}},n_{\text{bck}}$):
16:  while len(wpts) $\leq n_{\text{bck}}$ do
17:   $\mathbf{p}\leftarrow\texttt{wpts[-1]}$
18:   if $d(\mathbf{p},\mathbf{p}_{\text{goal}}^{\text{bck}})<\rho$ then
19:    break
20:   $\beta$ = getBearing($\mathbf{p},\mathbf{p}_{\text{goal}}^{\text{bck}}$)
21:   wpt $\leftarrow$ pointOnHeading($\mathbf{p},\beta,\rho_{\text{wpt}}$)
22:   if overshootsGoal($\texttt{wpt},\mathbf{p}_{\text{goal}}^{\text{bck}},\rho$) then
23:    wpt $\leftarrow\mathbf{p}_{\text{goal}}^{\text{bck}}$
24:   wpts.append(wpt)
25:  end while
As mentioned in Section III-A, several failures in transferring control instructions can occur in a row. Therefore, providing a single close waypoint is insufficient. From an operational standpoint, waypoint lists must extend well beyond a single dive so that the vehicle can continue its mission uninterrupted in the case of communication failures. The approach we propose uses the planner action to compute the first waypoint and several backup waypoints derived using straight-to-goal bearings. Since waypoint lists are recomputed and updated at each surfacing, this allows the vehicle to steer according to planner output during normal operation, and use a sensible backup trajectory otherwise.
Pseudocode for the waypoint list conversion procedure is given in Algorithm 3 and an example is shown in Figure 5. The algorithm takes as input the action $\alpha$, glider position $\mathbf{p}_{0}$, the position of the current goal $\mathbf{p}_{\text{goal}}^{\text{curr}}$, that of the next goal $\mathbf{p}_{\text{goal}}^{\text{next}}$, the goal tolerance radius $\rho$, the waypoint distance $\rho_{\text{wpt}}$, and number of backup waypoints $n_{\text{bck}}$.
@@ V-B Waypoint List Generation
The algorithm first determines the waypoint corresponding to the planner action (lines 3-12). If distance to the goal is low, the goal itself is used as the waypoint to avoid repeatedly oscillating around it. Otherwise, the waypoint is set along the heading of the planner action at the specified waypoint distance. Backup waypoints towards the current goal and, if required, the next goal (lines 13-14) are added.
The bearings for the backup waypoints are computed starting from the last waypoint (line 17). Additional backup waypoints towards a particular goal are not necessary if the previous waypoint is already within the same goal region (lines 18-19), and this serves as the loop termination condition. As mentioned, the bearing to the goal is used as the heading for the backup waypoints. If a backup waypoint overshoots the goal region, it is replaced with the goal itself (lines 22-23).
V-C Underwater Goal Achievement Detection
Regarding determining whether the goal was achieved, i.e., whether $\mathbf{s}\in\mathcal{G}$, there is an additional condition that must be checked in GLAE besides the dive surfacing within the goal region. The additional condition is necessary to align our problem formulation with operational characteristics of Slocum gliders, which detect whether a waypoint was achieved while underwater using dead reckoning (which we note may be inaccurate). As the waypoint matches the goal itself when approaching the goal region (see previous section), the glider may start automatically advancing towards the next goal while underwater.
Concretely, if the waypoint is deemed achieved underwater, the glider switches mid-dive to the next waypoint in the navigation plan. As the glider reports the waypoint it is following upon every surfacing, the position of the waypoint in the last navigation plan can be used to check if this condition is met. GLAE advances to the next goal in the ordered list if either this condition or the surfacing within the goal region condition is met.
V-D Instruction Set Generation
The second file type required, the instruction set, contains the dive control parameters $\mathbf{c}$. Following the guidance of glider pilots, GLAE supports specifying a number of fixed instruction sets $\mathcal{K}$ that determine the vehicle speed, sampling speed, and other operational variables. This ensures the control parameters that can be output by GLAE are within a field-tested, pre-approved range of configurations.
As previously mentioned, for simplicity, we assume that a fixed set of control parameters $\mathbf{c}$ is used for the duration of a particular mission. This assumption is realistic for the deployments considered in our evaluation, which require sampling of scientific data at regular time and distance intervals, and thus a fixed $\mathbf{c}$ is highly appropriate, as is also used in the field. A choice of a particular instruction set $\mathbf{c}\in\mathcal{K}$ corresponding, for example, to medium vehicle speed and high data sampling rates is specified when a mission is initialised.
V-E System Components and Integration
@@
Figure 6: Flow diagram illustrating the components and steps for online planning. This workflow is executed every few hours upon the surfacing of the glider.
Besides GLAE, several components are required to achieve the full control loop of the glider. This is made possible by integration with the Oceanids C2 infrastructure [37] built by the UK National Oceanography Centre (NOC). This platform provides the necessary components to achieve real-time ingestion of vehicle information and transfer of control instructions using standardised interfaces. In this section, we describe the system components and how they are integrated.

### v2

@@ 4.D Simulator Parameter Tuning
To enable tuning $\Theta$, we assume availability of a dataset $\mathcal{D}=\{\mathbf{x}^{(i)}\}_{i=1}^{N}$ of historical dives with entries $\mathbf{x}^{(i)}$ of the form $((\mathbf{s}^{(i)},\mathbf{a}^{(i)},\beta^{(i)},\hat{\mathbf{q}}^{(i)},\mathbf{b}^{(i)}),\mathbf{s}_{\text{true}}^{(i)})$, comprising the conditions prior to starting a dive and the true post-dive state after the surfacing of the glider. For simplicity, we also assume that all parameters $\mathbf{c}$ except $n_{\text{yos}}$ and $z_{\text{bottom}}$ are held constant within a dataset used to fit the simulator parameters.
@@
We focus on optimising the generated simulator positions and dive durations relative to the ground truths $\mathbf{p}^{(i)}_{\text{true}}$ and $\Delta\tau^{(i)}_{\text{true}}$ respectively, where $\Delta\tau=\tau^{\prime}-\tau$. For each $\mathbf{x}^{(i)}$, we draw $M$ samples of future states $\mathbf{s}^{(i,1)},\mathbf{s}^{(i,2)},\dots,\mathbf{s}^{(i,M)}\sim\mathcal{T}_{\Theta}(\mathbf{s},\mathbf{a},\beta,\hat{\mathbf{q}},\mathbf{b})$, recording the positions $\mathcal{P}^{(i)}=\{\mathbf{p}^{(i,1)},\mathbf{p}^{(i,2)},\dots,\mathbf{p}^{(i,M)}\}$ and the durations $\mathcal{I}^{(i)}=\{\Delta\tau^{(i,1)},\Delta\tau^{(i,2)},\dots,\Delta\tau^{(i,M)}\}$. We subsequently fit a per-dive Kernel Density Estimator (KDE) for simulator positions with a Gaussian kernel $\hat{p}^{(i)}_{\Theta}=\text{KDE}(\mathcal{P}^{(i)};h)$ and bandwidth $h$, and a Gaussian distribution $\widehat{\Delta\tau}^{(i)}_{\Theta}\sim\mathcal{N}(\mathcal{I}^{(i)};\mu^{(i)},\sigma^{2})$ for dive durations with a fixed $\sigma$.
Evaluating the probabilities $\hat{p}_{\Theta}^{(i)}(\mathbf{p}^{(i)}_{\text{true}})$ and $\widehat{\Delta\tau}_{\Theta}^{(i)}(\Delta\tau^{(i)}_{\text{true}})$ quantifies how well the simulator parameters $\Theta$ fit a particular dive. Furthermore, we also penalise the spread $\sum_{j=1}^{M}{(\mathbf{p}^{(i,j)}-\overline{\mathbf{p}^{(i,j)}})}^{2}$ of the positions relative to their mean. Without this penalty, the optimisation procedure tends to favour parameter configurations with overly diffuse transition distributions. Better scores are obtained on the worst dives as probability mass is placed on a wide range of outcomes. This behaviour is undesirable, as it leads to uninformative predictions that are also detrimental to the downstream planning. The scoring function $J(\Theta)$ is therefore defined as:
@@
$J(\Theta)=\sum_{i=1}^{N}{\log\hat{p}^{(i)}_{\Theta}(\mathbf{p}^{(i)}_{\text{true}})+\lambda_{\text{dur}}\sum_{i=1}^{N}{\log\widehat{\Delta\tau}^{(i)}_{\Theta}(\Delta\tau^{(i)}_{\text{true}})}}\\ -\lambda_{\text{reg}}\sum_{i=1}^{N}{\sum_{j=1}^{M}{(\mathbf{p}^{(i,j)}-\overline{\mathbf{p}^{(i,j)}})^{2}}},$
@@
where the first term corresponds to the log-likelihood $\ell(\mathcal{D};\Theta)$ of surfacing locations, the second term corresponds to the duration score with a configurable weight $\lambda_{\text{dur}}$, and the third term penalises the spread of samples generated for each $\mathbf{x}^{(i)}$ with a configurable weight $\lambda_{\text{reg}}$.
@@
We seek to find the optimal parameters $\Theta^{*}=\operatorname*{\arg\!\max}_{\Theta}(J(\Theta))$. This high-dimensional optimisation problem has a scoring function that is expensive to evaluate and cannot be computed in closed form. As such, Bayesian optimization is a suitable framework. We opt for BoTorch [32, 33], whose acquisition function navigates the exploration-exploitation tradeoff to generate promising candidate parameter sets $\Theta_{k}$ at each iteration $k$ of a total $n_{\text{SIM-BO}}$ iterations.
4.E Sample-Based Online Planner
@@
7:  $\mathbf{a}\leftarrow$ MajorityVote(allActions)
8:  return $\mathbf{a}$
9:Function RunUCT(II, randomState):
10:  $(\mathbf{p}_{0},\tau_{0},\mathbf{p}_{\text{goal}},\rho,\mathbf{q}(\cdot),\mathbf{b}(\cdot))\leftarrow\ I$
11:  create root node $v_{\text{root}}$ from $s_{0}=(\mathbf{p}_{0},\tau_{0})$
12:  for $i$ in $1,n_{\text{trials}}]$ do
13:   $v_{\text{leaf}}\leftarrow$ TreePolicy($v_{\text{root}}$)
14:   $\mathbf{s}\leftarrow$ GetState($v_{\text{leaf}}$)
15:   if $\mathbf{s}\in\mathcal{G}$ then
16:    $r\leftarrow$ SumCosts($\mathbf{s}_{0},\mathbf{s}$)
17:   else
18:    $r\leftarrow\epsilon_{\text{heur}}*(d(\mathbf{p},\mathbf{p}_{\text{goal}})/s)$
19:   Backup($v_{\text{leaf}},r$)
20:  action $\leftarrow$ RobustChild($v_{\text{root}}$)
21:  return action
Gliders surface every few hours and await instructions for the next dive while uploading scientific data. Thus, an online planning approach is natural for solving the MDP presented in Section 4.B, leveraging the time on the surface to derive a navigation plan. At each surfacing, the problem instance $I=(\mathbf{p}_{0},\tau_{0},\mathbf{p}_{\text{goal}},\rho,\mathbf{q}(\cdot),\mathbf{b}(\cdot))$ is assembled using the position $\mathbf{p}_{0}$ from the glider’s GPS fix, time $\tau_{0}$, goal specification, current forecasts, and bathymetry.
To perform sample-based online planning and solve the MDP approximately, we leverage Monte Carlo Tree Search (MCTS), an online search algorithm for determining the best action to apply at the current state of an MDP. As a full review of MCTS is out of scope, we refer the interested reader to [26] for more details. Specifically, we opt for the Upper Confidence Bounds for Trees (UCT) [34] variant, which treats action selection at each node as an independent multi-armed bandit problem, applying the principles behind the Upper Confidence Bounds (UCB) [35] algorithm to the tree search setting. Planning is carried out while the glider is at the surface using $n_{\text{trials}}$ search tree traversals. $n_{\text{trials}}$ is a parameter that is set on a per-deployment basis to account for the expected time duration of the glider at the surface and the computational resources available.
Pseudocode for the planner is given in Algorithm 2. We use the following enhancements over the basic MCTS scheme. Our planner leverages double progressive widening, which is suitable for scenarios with stochastic transitions and continuous state spaces [36, 37]. It is able to achieve a suitable compromise between bias (induced by the first expanded nodes) and variance (in the result of trials for each state). We note that, despite the discretisation of the actions, we opt for progressive widening in the action space also. This is done to keep the branching factor low and increase the tree depth, enabling us to plan over a longer horizon.
Furthermore, we use root parallelism in order to take advantage of multi-core architectures despite the sequential nature of MCTS. This creates $n_{\text{threads}}$ search trees that progress independently from the given root state. The final results are aggregated by majority voting to obtain a single action [38, 39]. This improves the robustness of the navigation plan to both noise in the current forecasts and action choices early in the tree search.
We also highlight the following differences from a standard implementation:
  • Use of heuristic. We do not perform rollouts to estimate the value of leaf nodes as sampling transitions requires running a computationally intensive simulator. This is impractical given the time budget available at each surfacing. Instead, we use a heuristic (line 18) that estimates the time to reach the goal based on the distance to goal and constant speed. The error term $\epsilon_{\text{heur}}$ of the heuristic was calibrated based on preliminary deployment data and was set to approximately $1.77$. Equivalently, the typical trajectory length was $1.77\times$ that of the straight line.
5 GLIDER AUTONOMY LONG-TERM PLANNING ENGINE
@@
In this section, we describe the components required for translating our methodological and algorithmic contributions discussed in the previous section to a field-validated system that performs long-term autonomous navigation planning for underwater gliders. The Glider Autonomy Long-term Planning Engine was designed and implemented in collaboration with operational stakeholders in order to enhance operational practicality and safety.
5.A Automated Goal Management
@@
The previous section specified problem instances $I$ with a single goal $\mathbf{p}_{\text{goal}}$. However, glider missions typically consist of many such goals. A common use case, which we target, is to have the glider sample data between a set of predefined points, i.e., a transect in the case of two points, and the corners of a polygon more generally. We therefore assume we are given an ordered list of points $\{\mathbf{p}^{(i)}_{\text{goal}}\}_{i=1}^{n_{\text{goals}}}$. The system advances to the next goal in the list upon achieving the current one, and cycles to the beginning of the list once the end is reached.
For determining whether the goal was achieved, i.e., whether $\mathbf{s}\in\mathcal{G}$, an additional condition must be checked in GALE besides surfacing within the goal region. A goal may also be detected as having been achieved while underwater using dead reckoning (which we note may be inaccurate). This reflects the operation of the on-board glider software.
@@
Figure 8: Diagram illustrating the components and steps of the two system workflows. The Mission Configuration Workflow (right) is executed once at the start the mission, while the Online Planning Workflow (left) is executed every few hours upon the surfacing of the glider.
5.B Translation of Control Instructions
Actions $\mathbf{a}=(\alpha,\mathbf{c})$ output by the planner must be translated to files that can be understood by a Slocum glider. Two types of files are required: a waypoint list specifying an ordered sequence of coordinates that the glider should reach, and an instruction set containing the control parameters $\mathbf{c}$.
To generate the waypoint list, we use the $\alpha$ component of the planner action for computing the first waypoint and several backup waypoints derived using straight-to-goal bearings. Since waypoint lists are recomputed and updated at each surfacing, the vehicle can steer according to planner output during normal operation and use a sensible backup trajectory otherwise. As mentioned in Section 2, several failures in transferring control instructions can occur in a row. Therefore, waypoint lists must extend well beyond a single dive so that the vehicle can continue its mission uninterrupted in the case of communication failures.
A waypoint list generation example is shown in Figure 7. The algorithm, which we describe at a high level due to space limitations, takes as input the action $\alpha$, glider position $\mathbf{p}_{0}$, the position of the current goal $\mathbf{p}_{\text{goal}}^{\text{curr}}$, the position of the next goal $\mathbf{p}_{\text{goal}}^{\text{next}}$, the goal tolerance radius $\rho$, the waypoint distance $\rho_{\text{wpt}}$, and number of backup waypoints $n_{\text{bck}}$. The algorithm first determines the waypoint corresponding to the planner action. If distance to the goal is low, the goal itself is used as the waypoint to avoid oscillating around it. Otherwise, the waypoint is set along the heading of the planner action at the specified waypoint distance. Backup waypoints towards the current goal and, if required, the next goal are added. The bearings for the backup waypoints are computed starting from the previous waypoint. Additional backup waypoints towards a particular goal are not necessary if the previous waypoint is already within the same goal region. If a backup waypoint overshoots the goal region, it is replaced with the goal itself.
Generating the instruction set file involves a format translation of the control parameters $\mathbf{c}$. Following the guidance of glider pilots, GALE supports specifying a number of fixed instruction sets $\mathcal{K}$ from which $\mathbf{c}$ is chosen when the mission is initialised. This ensures the control parameters are within a field-tested range of configurations, protecting the assets and ensuring pilot trust in the system. As previously mentioned, for simplicity, we assume that a fixed $\mathbf{c}$ is used for the duration of a particular mission. This assumption is realistic for the deployments considered in our evaluation, which require sampling of scientific data at regular time and distance intervals. Dynamically varying the instruction set on a per-action basis is left for future work.
5.C System Components and Integration